# Perception Stitching: Zero-Shot Perception Encoder Transfer for Visuomotor Robot Policies

### v1

@@ 4.1 Simulation Experiment Setup and Baselines
Push Devin et al. 2017 60.7$\pm$10.6 8.7$\pm$4.99 16.7$\pm$3.77 59.3$\pm$6.80 29.3$\pm$7.36 19.3$\pm$5.73 32.3
Cannistraci et al. 2024 (linear) 89.3$\pm$4.11 94.0$\pm$2.83 64.7$\pm$1.89 74.7$\pm$6.18 74.0$\pm$2.83 78.7$\pm$2.49 79.2
Cannistraci et al. 2024 (non-linear) 12.7$\pm$1.89 18.7$\pm$4.99 42.8$\pm$3.27 23.3$\pm$0.94 6.0$\pm$4.32 5.3$\pm$2.49 18.1
PeS (w/o disent. loss) 100.0$\pm$0.0 86.0$\pm$2.83 80.7$\pm$9.84 100.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0 94.5
PeS (w. l1 & l2 loss) 88.7$\pm$4.99 95.3$\pm$1.89 90.0$\pm$5.66 100.0$\pm$0.0 93.3$\pm$0.94 80.7$\pm$4.99 91.3
PeS 100.0$\pm$0.00 100.0$\pm$0.00 95.3$\pm$0.94 100.0$\pm$0.0 92.7$\pm$2.50 100.0$\pm$0.00 98
Lift Devin et al. 2017 0.0$\pm$0.00 5.3$\pm$2.49 48.0$\pm$5.89 9.3$\pm$4.11 14.7$\pm$4.99 36.0$\pm$1.63 18.9
Cannistraci et al. 2024 (linear) 72.7$\pm$3.77 64.0$\pm$2.83 86.0$\pm$4.32 68.7$\pm$1.88 88.7$\pm$1.88 57.3$\pm$2.49 72.9
Cannistraci et al. 2024 (non-linear) 89.3$\pm$2.49 36.0$\pm$3.27 52.7$\pm$3.40 93.3$\pm$2.49 16.7$\pm$2.49 21.3$\pm$0.94 51.6
PeS (w/o disent. loss) 83.3$\pm$6.60 80.7$\pm$5.73 93.3$\pm$0.94 91.3$\pm$5.73 79.3$\pm$2.49 93.3$\pm$2.49 86.9
PeS (w. l1 & l2 loss) 97.3$\pm$2.49 85.3$\pm$0.94 90.7$\pm$0.94 86.0$\pm$4.32 88.0$\pm$1.63 84.7$\pm$3.77 88.7
PeS 92.7$\pm$2.50 94.7$\pm$1.89 89.3$\pm$4.11 96.0$\pm$1.63 88.7$\pm$0.94 93.0$\pm$0.03 92.4
Table 1: Zero-Shot Transfer Success Rates in basic Simulation tasks.
Mask Zoom in Blurred Noise Fisheye Camera Position Average
Can Devin et al. 2017 19.3$\pm$5.25 24.7$\pm$1.89 2.67$\pm$1.89 6.0$\pm$4.32 29.3$\pm$3.40 1.3$\pm$1.89 13.9
Cannistraci et al. 2024 (linear) 33.3$\pm$0.94 48.0$\pm$1.63 48.7$\pm$2.49 65.3$\pm$0.94 26.7$\pm$3.77 34.7$\pm$3.77 42.8
Cannistraci et al. 2024 (non-linear) 72.7$\pm$0.94 24.7$\pm$2.49 37.3$\pm$4.99 42.7$\pm$3.40 8.7$\pm$1.89 39.3$\pm$1.89 37.6
PeS (w/o disent. loss) 44.7$\pm$8.06 89.3$\pm$4.11 34.7$\pm$4.11 30.7$\pm$6.80 92.7$\pm$2.50 44.7$\pm$3.40 56.1
PeS (w. l1 & l2 loss) 47.3$\pm$0.94 58.7$\pm$1.88 54.0$\pm$8.64 36.0$\pm$7.12 58.7$\pm$1.88 64.7$\pm$6.60 53.2
PeS 83.3$\pm$5.24 89.3$\pm$2.49 74.0$\pm$2.83 78.7$\pm$4.11 56.0$\pm$2.83 78.7$\pm$2.49 76.7
Stack Devin et al. 2017 0.7$\pm$0.94 8.0$\pm$1.63 0.7$\pm$0.94 24.0$\pm$2.83 0.0$\pm$0.00 14.0$\pm$3.27 7.9
Cannistraci et al. 2024 (linear) 47.3$\pm$0.94 62.0$\pm$4.32 32.7$\pm$3.77 30.7$\pm$0.94 54.0$\pm$8.64 14.7$\pm$6.18 40.2
Cannistraci et al. 2024 (non-linear) 10.0$\pm$1.63 12.0$\pm$0.00 0.0$\pm$0.00 3.3$\pm$0.94 0.0$\pm$0.00 0.7$\pm$0.94 4.3
PeS (w/o disent. loss) 34.0$\pm$11.43 10.7$\pm$4.11 62.0$\pm$10.71 34.0$\pm$7.12 22.7$\pm$3.77 26.0$\pm$4.32 31.6
PeS (w. l1 & l2 loss) 92.7$\pm$0.94 98.0$\pm$0.00 62.7$\pm$6.60 24.0$\pm$4.90 59.3$\pm$7.36 58.7$\pm$1.88 65.9
PeS 94.7$\pm$0.94 96.7$\pm$0.94 90.0$\pm$1.63 96.7$\pm$1.89 97.3$\pm$2.49 80.0$\pm$4.90 92.6
Door Open Devin et al. 2017 9.3$\pm$4.11 5.3$\pm$0.94 0.0$\pm$0.00 4.0$\pm$1.63 0.7$\pm$0.94 0.0$\pm$0.00 3.2
Cannistraci et al. 2024 (linear) 0.0$\pm$0.00 1.3$\pm$0.94 10.7$\pm$2.49 10.7$\pm$4.99 2.0$\pm$1.63 47.3$\pm$9.29 12
Cannistraci et al. 2024 (non-linear) 26.0$\pm$2.83 31.3$\pm$4.99 49.3$\pm$8.22 48.0$\pm$5.89 62.7$\pm$3.40 44.7$\pm$3.40 43.7
PeS (w/o disent. loss) 24.7$\pm$7.71 44.0$\pm$2.83 34.7$\pm$3.77 0.7$\pm$0.94 36.7$\pm$0.94 23.3$\pm$3.40 27.4
PeS (w. l1 & l2 loss) 4.0$\pm$1.63 78.0$\pm$5.66 3.3$\pm$0.94 2.0$\pm$1.63 42.7$\pm$4.99 6.0$\pm$3.26 22.7
PeS 58.7$\pm$4.11 68.7$\pm$0.94 70.7$\pm$0.94 52.7$\pm$3.40 64.7$\pm$4.99 48.7$\pm$3.40 60.7
Table 2: Zero-Shot Transfer Success Rates in difficult Simulation tasks.
Tab. 2 reports the performance in more difficult tasks: Can, Stack, and Door Open. PeS achieves 60% to 93% of success rates, while the success rates of all three baselines are lower than 43%. This result suggests that previous methods cannot solve the zero-shot transfer problem for difficult tasks with satisfactory performance, which usually involve complicated visual background (e.g., Can), long-horizon multi-stage motions (e.g., Stack), and articulated objects (e.g., Door Open). We also notice that the two ablation methods have significant performance drops (20% to 60%) in these difficult tasks, showing that the disentanglement regularization in PeS largely improves the zero-shot transfer performance. Although using L1 and L2 regularization also leads to good success rates in some cases, the disentanglement regularization has more consistent performance and higher success rates on average.
@@ 4.2 Real World Experiment
Camera Configurations We conducted Reach, Push, Lift, and Stack in our real-world experiments (Fig. 5). In Reach, the robot starts from some random position in the air and reaches the red cube. In this experiment, the policy $1$ is trained with a front-view camera with a broken lens (Fig. 5) and a normal in-hand camera. The policy $2$ is trained with a normal front-view camera and an in-hand camera with a broken lens. Then we stitch the front view encoder of the policy $2$ to the in-hand view encoder and action decoder of the policy $1$ and test the stitched policy with two normal cameras. In Push, the robot starts from some random position to reach the cube and then pushes the cube across the green line on the table. The training and testing process of Push is similar to that of the Reach, and the only difference is replacing the camera with a broken lens with a camera with a masked lens, as shown in Fig. 5. In Lift, the red cube is placed at a random position on the table with a random orientation. The robot needs to grab the cube and lift it up. In this experiment, the policy $1$ is trained with in-hand and side-view cameras. The policy $2$ is trained with side-view and front-view cameras. Then we stitch the front view encoder of the policy $2$ to the in-hand view encoder and action decoder of the policy $1$, and test the stitched policy with an in-hand camera and a front-view camera, as shown in the left half part of the Fig. 2. In Stack, the red cube is placed at a random position on the left side of the table, and the robot should lift it up and stack it on the blue cube on the right. Similar to Lift, it also involves the stitching between cameras at different positions (left, front, and right). Stack is the most difficult one in all the real-world experiments. It requires the most sophisticated manipulation skills and long-horizon multi-stage motions.
@@
Table 3: Zero-Shot Transfer Success Rates in Real World.
As shown in Tab. 3, PeS achieves 100% of success rate in Reach, 85% in Push, 80% in Lift, and 45% in Stack, while Devin et al. 2017 baseline gets 0% of success rates in all the experiments. In the experiments, we observe that PeS policies show more accurate motions compared with the baseline policies. In comparison, Devin et al. 2017 baseline policies cannot output actions accurately enough to accomplish the tasks, although they have some attempts to complete the tasks in some cases. Take the Lift task as an example, the robot trained with the baseline method reaches some positions close to the cube but always fails to grasp it. In the most difficult Stack task, PeS robot can still achieves many success cases, while the baseline robot cannot output meaningful motions. Please refer to the supplementary material for the experiment videos.
In summary, we find that the performance advantage of PeS is more pronounced in the real world than that in the simulation. We assume that this is because the real-world images have more complicated components than the simulation images, and there are more noises and disturbances in the real world. To the best of our knowledge, PeS is the first method that enables vision-based zero-shot transfer in the real world via reassembling neural network components.
@@ 4.3 Analysis: Latent Space at Module Interface
Figure 6: Latent Space Analysis. (a) We visualize the latent representations of the front-view encoder of policy $1$ and the side-view encoder of policy $2$ of the Push-Camera Position experiment. The $256$D representations are reduced to $3$D with PCA [30]. The red dots represent samples where the robot’s end effector is at higher positions, blue dots indicate medium heights and green dots correspond to lower positions near the cube. We compare our (PeS) method with the baseline (Devin et al. 2017) method. (b) & (c) We compare the distances of the latent representations in all the experiments in the Push task. One representation is from the second view encoder of policy $1$ and the other is from the second view encoder of policy $2$.)
@@
We choose the Push-Camera Position experiment in the simulation and visualize the corresponding latent representations to have an intuitive understanding of the mechanism of PeS. We first reduced the $256$D representations to $3$D with PCA[30] for visualization. Since the side view encoder of the policy $2$ is stitched to the policy $1$ at the position of its original front view encoder, we compare the latent representations of these two encoders. As shown in Fig. 6(a), the latent representations trained with PeS have similar shapes to each other. In comparison, the latent representations with the Devin et al. 2017 baseline are not similar but have an approximately isometric transformation (rotation in this case) relationship with each other.
For further quantitative analysis, we select the Push task in the simulation and calculate the pairwise distances of the latent representations in all Push experiments with different visual configurations. Fig. 6 (b) shows that PeS significantly reduces the cosine distances of the latent representations in all these experiments. Fig. 6 (c) shows that the L2 distances with PeS are generally smaller than that with the Devin et al. 2017 baseline, but the differences are not distinguishable in some cases (e.g. Push-Zoom in). This observation is different from Jian et al. [38]’s work where they embed the low dimensional observations with the task modules and find that both cosine and L2 distances of the relative representations are significantly smaller.
@@ 4.4 Analysis: Highlight Attention Regions with Grad-CAM
We visualize the attention map from the policy modules to further explain why certain policies work well while others fail. We adapt Grad-CAM [74] to help with this visualization.
@@
We choose the Lift-Camera Position experiment and visualize the attention heatmap of the stitched policies with PeS and the baseline separately. As shown in Fig. LABEL:fig:inhand_heatmap, from the in-hand camera, our policy is focusing on the cube between the two gripper fingers, which is intuitively what humans pay attention to in a Lift task. In comparison, the Devin et al. 2017 baseline policy has more attention to the upper middle part of the image, which is the desk surface and is not informative for this task. In Fig. LABEL:fig:side_heatmap, our policy is paying attention to the robot end effector from the side-view camera, while the baseline policy has more attention to the two sides of the image, while these regions have no crucial object. To sum up, the attention maps from the Grad-CAM suggest that the stitched policy with PeS performs better than the baseline because it can pay attention to the crucial regions for accomplishing the task, while the baseline policy cannot. This result provides an intuitive explanation of the good performance of PeS. Videos of the attention heatmaps during the manipulation process can be found in the supplementary material, and we provide the mathematical details of the Grad-CAM in Appendix H.
@@ 5 Discussion
Conclusion We present Perception Stitching (PeS), a method for zero-shot visuomotor policies transfer via latent spaces alignment. PeS aligns the latent spaces of different visual encoders by enforcing the relative representations invariant to isometric transformations and, therefore, allows the trained visual encoders to be reused in a plug-and-go manner. Our evaluation covers $30$ simulation experiments and $4$ real-world experiments with a variety of manipulation tasks and camera configurations. The results demonstrate significant performance improvement with PeS for zero-shot transfer, and its advantage is especially pronounced in real-world tasks. Moreover, we conduct quantitative and qualitative analysis to understand the mechanism of the superior performance of PeS. We hope that this work can inspire further exploration of the compositionality and modularity in robot transfer learning.
Limitations and Future Work Though there is no additional training required during transfer, in order to obtain the anchor states, our framework still requires the robot to replay the trajectories in the previous dataset to collect a new dataset and then select anchors with the corresponding indices. Although this is feasible, as we have demonstrated in our real-world experiments, the trajectories replaying process in the real world usually takes longer time than collecting a new dataset by random sampling. Future work can develop an alternative algorithm for more efficient anchor selection. Another promising direction is to explore the combination of PeS with other robot transfer learning techniques that focus beyond vision encoder transfer, such as embodiment transfer and scale to large robot learning datasets.
Acknowledgments and Disclosure of Funding
This work is supported by ARL STRONG program under awards W911NF2320182 and W911NF2220113, by DARPA FoundSci program under award HR00112490372, and by AFOSR under award #FA9550-19-1-0169.
@@ Appendix B Perception Stitching Algorithm
Optimize $f^{E_{1}}(g^{E_{1}}_{1}(o^{E_{1}}_{1}),g^{E_{1}}_{2}(o^{E_{1}}_{2}),o^{E_{1}}_{l},\mathbb{A}_{1})$ with dataset $\mathbb{D}_{1}$ with the Behavior Cloning algorithm [73].
Optimize $f^{E_{2}}(g^{E_{2}}_{1}(o^{E_{2}}_{1}),g^{E_{2}}_{2}(o^{E_{2}}_{2}),o^{E_{2}}_{l},\mathbb{A}_{2})$ with dataset $\mathbb{D}_{2}$ with the Behavior Cloning algorithm [73].
@@
The algorithm of the Perception Stitching (PeS) is shown in Algorithm 1. The first step of this algorithm is to collect an expert dataset $\mathbb{D}_{1}$ in the environment $E_{1}$ by teleoperating the robot by a proficient human expert. Then we replay the expert trajectories in $\mathbb{D}_{1}$ to collect another expert dataset $\mathbb{D}_{2}$ in the environment $E_{2}$. We collect the anchor set $\mathbb{A}_{1}$ in the dataset $\mathbb{D}_{1}$ via the K-means algorithm [24]. Then we collect the anchor set $\mathbb{A}_{2}$ which has the same indices in $\mathbb{D}_{2}$ as $\mathbb{A}_{1}$ in $\mathbb{D}_{1}$. For the environment $E_{3}$ with two visual configurations $o_{1}^{E_{1}}$ and $o_{2}^{E_{2}}$, we assemble an anchor set $\mathbb{A}_{3}=\{a_{3}^{(j)}\}$ consists of images of $o_{1}^{E_{1}}$ from $\mathbb{A}_{1}$ and images of $o_{2}^{E_{2}}$ from $\mathbb{A}_{2}$. In the next step, we train the two polices $f^{E_{1}}(g^{E_{1}}_{1}(o^{E_{1}}_{1}),g^{E_{1}}_{2}(o^{E_{1}}_{2}),o^{E_{1}}_{l},\mathbb{A}_{1})$ and $f^{E_{2}}(g^{E_{2}}_{1}(o^{E_{2}}_{1}),g^{E_{2}}_{2}(o^{E_{2}}_{2}),o^{E_{2}}_{l},\mathbb{A}_{2})$ in $E_{1}$ and $E_{2}$ with the Behavior Cloning algorithm [73] separately. Then we initialize a stitched policy $f^{E_{1}}(g^{E_{1}}_{1}(o^{E_{1}}_{1}),g^{E_{2}}_{2}(o^{E_{2}}_{2}),o^{E_{1}}_{l},\mathbb{A}_{3})$. This stitched policy is tested in the task $T$ in $E_{3}$. Its performance is measured by its success rate.

### v2

@@ 4.2 Simulation Experiments with Single Camera View and Triple Camera Views
Triple Camera Views We pick the Stack task and train policies with three visual encoders that embed three different visual configurations. For the first five visual configurations that add different effects to the observations, policies are all trained with the in-hand view, close-up view, and side view. Policy 1 has the effect (e.g. Gaussian noise) only in vision 3, and policy 2 has the same effect in vision 1 and vision 2. Then we stitch the encoder 3 of policy 2 to the rest parts of policy 1 and test the stitched policy with three normal cameras. For the last experiment that involves different camera positions, policy 1 is trained with the in-hand view, close-up view, and front view, and policy 2 is trained with close-up view, front view, and side view. Then we stitch the side view encoder of policy 2 to the in-hand view, close-up view encoders, and decoder of policy 1. The stitched policy is tested with in-hand view, close-up view, and side view. This triple camera perception stitching process is presented in Figure 6.
@@ 4.3 Real World Experiment
Camera Configurations We conducted Reach, Push, Lift, and Stack in our real-world experiments (Fig. 7). In Reach, the robot starts from some random position in the air and reaches the red cube. In this experiment, the policy $1$ is trained with a front-view camera with a broken lens (Fig. 7) and a normal in-hand camera. The policy $2$ is trained with a normal front-view camera and an in-hand camera with a broken lens. Then we stitch the front view encoder of the policy $2$ to the in-hand view encoder and action decoder of the policy $1$ and test the stitched policy with two normal cameras. In Push, the robot starts from some random position to reach the cube and then pushes the cube across the green line on the table. The training and testing process of Push is similar to that of Reach, and the only difference is replacing the camera with a broken lens with a camera with a masked lens, as shown in Fig. 7. In Lift, the red cube is placed at a random position on the table with a random orientation. The robot needs to grab the cube and lift it up. In this experiment, the policy $1$ is trained with in-hand and side-view cameras. The policy $2$ is trained with side-view and front-view cameras. Then we stitch the front view encoder of the policy $2$ to the in-hand view encoder and action decoder of the policy $1$, and test the stitched policy with an in-hand camera and a front-view camera, as shown in the left half part of the Fig. 2. In Stack, the red cube is placed at a random position on the left side of the table, and the robot should lift it up and stack it on the blue cube on the right. Similar to Lift, it also involves the stitching between cameras at different positions (left, front, and right). Stack is the most difficult one in all the real-world experiments. It requires the most sophisticated manipulation skills and long-horizon multi-stage motions.
@@
Table 5: Zero-Shot Transfer Success Rates in Real World. We report the success rates of three manipulation tasks in real world. Each success rate is calculated with $20$ games. PeS achieves a 100% success rate in the easiest Reach task, 85% in the Push task, 80% in the Lift task, and 45% in the hardest Stack task. In comparison, the baseline method doesn’t have any success case in all these tasks.
Results As shown in Tab. 5, PeS achieves 100% of success rate in Reach, 85% in Push, 80% in Lift, and 45% in Stack, while Devin et al. 2017 baseline gets 0% of success rates in all the experiments. In the experiments, we observe that PeS policies show more accurate motions compared with the baseline policies. In comparison, Devin et al. 2017 baseline policies cannot output actions accurately enough to accomplish the tasks, although they have some attempts to complete the tasks in some cases. Take the Lift task as an example, the robot trained with the baseline method reaches some positions close to the cube but always fails to grasp it. In the most difficult Stack task, PeS robot can still achieve many success cases, while the baseline robot cannot output meaningful motions. Please refer to the supplementary material for the experiment videos.
In summary, we find that the performance advantage of PeS is more pronounced in the real world than in the simulation. We assume that this is because there are more noises and disturbances in the real world. To the best of our knowledge, PeS is the first method that enables vision-based zero-shot transfer in the real world via reassembling neural network components.
@@ 4.4 Analysis: Latent Space at Module Interface
We choose the Push-Camera Position experiment in the simulation and visualize the corresponding latent representations to have an intuitive understanding of the mechanism of PeS. We first reduced the $256$D representations to $3$D with PCA(Hotelling, 1933) for visualization. Since the side view encoder of the policy $2$ is stitched to the policy $1$ at the position of its original front view encoder, we compare the latent representations of these two encoders. As shown in Fig. 8(a), the latent representations trained with PeS have similar shapes to each other. In comparison, the latent representations with the Devin et al. 2017 baseline are not similar but have an approximately isometric transformation (rotation in this case) relationship with each other. The latent representations visualization of other tasks can be found in Appendix F.
Figure 8: Latent Space Analysis. (a) We visualize the latent representations of the front-view encoder of policy $1$ and the side-view encoder of policy $2$ of the Push-Camera Position experiment. The $256$D representations are reduced to $3$D with PCA (Hotelling, 1933). The red dots represent samples where the robot’s end effector is at higher positions, blue dots indicate medium heights and green dots correspond to lower positions near the cube. We compare our (PeS) method with the baseline (Devin et al. 2017) method. (b) & (c) We compare the distances of the latent representations in all the experiments in the Push task. One representation is from the second view encoder of policy $1$ and the other is from the second view encoder of policy $2$.)
For further quantitative analysis, we select the Push task in the simulation and calculate the pairwise distances of the latent representations in all Push experiments with different visual configurations. Fig. 8 (b) shows that PeS significantly reduces the cosine distances of the latent representations in all these experiments. Fig. 8 (c) shows that the L2 distances with PeS are generally smaller than that with the Devin et al. 2017 baseline, but the differences are not distinguishable in some cases (e.g. Push-Zoom in). More mathematical details of calculating the cosine and L2 distances can be found in Appendix C.
@@ 4.5 Analysis: Highlight Attention Regions with Grad-CAM
We visualize the attention map from the policy modules to further explain why certain policies work well while others fail. To this end, we modify the Gradient-weighted Class Activation Mapping (Grad-CAM) (Selvaraju et al., 2017) approach to highlight the regions that the policies pay attention to. Grad-CAM is widely applied to neural classification models with convolutional layers. To adapt the Grad-CAM from image classification to robot learning, we replace the before-softmax score $y^{c}$ for class $c$ of the image classification networks with the log-likelihood $l(a)$ of the robot action $a$ in the training dataset. We denote the $k^{th}$ feature map activation output from the last convolutional layer as $A^{k}$. Then the backpropagated gradient of $l(a)$ with respect to $A^{k}$ is computed as $\frac{\partial l(a)}{\partial A^{k}}$. We do global average pooling of these gradients over the width (indexed by $i$) and height (indexed by $j$) dimensions of the feature map to get the neuron importance weight :
This weight $\alpha^{a}_{k}$ captures the "importance" of feature map $k$ for robot action $a$. Then, the attention map $L_{\text{Grad-CAM }}^{a}$ is calculated as the weighted combination of forward activation maps followed by a ReLU (Selvaraju et al., 2017):
We apply ReLU because we are only interested in the features that have a positive influence on the actions. The intensity of these pixels should be increased in order to increase the log-likelihood $l(a)$ (Selvaraju et al., 2017). This $L_{\text{Grad-CAM }}^{a}$ is a heatmap of the same size as the convolutional feature maps $A^{k}$. We upsample it to the input image size with bilinear interpolation to get the final attention heatmap of the input image (Selvaraju et al., 2017). A larger value on this heatmap means this pixel contributes to a larger gradient of the log-likelihood of the robot action.
We choose the Lift-Camera Position experiment and visualize the attention heatmap of the stitched policies with PeS and the baseline separately. As shown in Fig. LABEL:fig:inhand_heatmap, from the in-hand camera, our policy is focusing on the cube between the two gripper fingers, which is intuitively what humans pay attention to in a Lift task. In comparison, the Devin et al. 2017 baseline policy has more attention to the upper middle part of the image, which is the desk surface, and is not informative for this task. In Fig. LABEL:fig:side_heatmap, our policy is paying attention to the robot end effector from the side-view camera, while the baseline policy has more attention to the two sides of the image, while these regions have no crucial object. To sum up, the attention maps from the Grad-CAM suggest that the stitched policy with PeS performs better than the baseline because it can pay attention to the crucial regions for accomplishing the task, while the baseline policy cannot. This result provides an intuitive explanation of the good performance of PeS. Videos of the attention heatmaps during the manipulation process can be found in the supplementary material.
@@ 5 Discussion
Conclusion We present Perception Stitching (PeS), a method for zero-shot visuomotor policies transfer via latent spaces alignment. PeS aligns the latent spaces of different visual encoders by enforcing the relative representations invariant to isometric transformations and, therefore, allows the trained visual encoders to be reused in a plug-and-go manner. Our evaluation covers $35$ simulation experiments and $4$ real-world experiments with a variety of manipulation tasks and camera configurations. The results demonstrate significant performance improvement with PeS for zero-shot transfer, and its advantage is especially pronounced in real-world tasks. Moreover, we conduct quantitative and qualitative analyses to understand the mechanism of the superior performance of PeS. We hope that this work can inspire further exploration of the compositionality and modularity in robot learning.
Limitations and Future Work There is no additional training required during transfer, but to obtain the anchor states our framework still requires the robot to replay the trajectories in the previous dataset to collect a new dataset and then select anchors with the corresponding indices. Although this is feasible, as we have demonstrated in our real-world experiments, the trajectories replaying process in the real world usually takes longer time than collecting a new dataset by random sampling. Future work can develop an alternative algorithm for more efficient anchor selection. We also acknowledge that PeS cannot yet achieve perfect performance on some tasks that require very high precision or have very long horizons. One major cause of the potential failure is that the PeS can only enforce an approximate invariance of the latent representations but not a strict invariance, which leads to the loss of some accuracy of the stitched policy. It is an interesting topic for future work to further refine the latent space alignment.
In addition, the policy trained with the Behavior Cloning algorithm usually cannot recover from the error accumulation very well in long-horizon tasks. Although our paper focuses on imitation learning, one exciting future direction is to explore our techniques in a reinforcement learning setup (Ricciardi et al., 2024) or explore more advanced imitation learning that handles error accumulation, since these approaches have shown some capabilities of recovery from failures in recent works. Another promising direction is to explore the combination of PeS with other robot transfer learning techniques that focus beyond vision encoder transfer, such as embodiment transfer and scale to large robot learning datasets.
Acknowledgement
This work is supported by ARL STRONG program under awards W911NF2320182 and W911NF2220113, by DARPA FoundSci program under award HR00112490372, by DARPA TIAMAT HR00112490419, and by AFOSR under award #FA9550-19-1-0169.
@@ Appendix B Perception Stitching Algorithm
Optimize $f^{E_{1}}(g^{E_{1}}_{1}(o^{E_{1}}_{1}),g^{E_{1}}_{2}(o^{E_{1}}_{2}),o^{E_{1}}_{l},\mathbb{A}_{1})$ with dataset $\mathbb{D}_{1}$ with the Behavior Cloning algorithm Schaal (1996).
Optimize $f^{E_{2}}(g^{E_{2}}_{1}(o^{E_{2}}_{1}),g^{E_{2}}_{2}(o^{E_{2}}_{2}),o^{E_{2}}_{l},\mathbb{A}_{2})$ with dataset $\mathbb{D}_{2}$ with the Behavior Cloning algorithm Schaal (1996).
@@
The algorithm of the Perception Stitching (PeS) is shown in Algorithm 1. The first step of this algorithm is to collect an expert dataset $\mathbb{D}_{1}$ in the environment $E_{1}$ by teleoperating the robot by a proficient human expert. Then we replay the expert trajectories in $\mathbb{D}_{1}$ to collect another expert dataset $\mathbb{D}_{2}$ in the environment $E_{2}$. We collect the anchor set $\mathbb{A}_{1}$ in the dataset $\mathbb{D}_{1}$ via the K-means algorithm Hartigan & Wong (1979). Then we collect the anchor set $\mathbb{A}_{2}$ which has the same indices in $\mathbb{D}_{2}$ as $\mathbb{A}_{1}$ in $\mathbb{D}_{1}$. For the environment $E_{3}$ with two visual configurations $o_{1}^{E_{1}}$ and $o_{2}^{E_{2}}$, we assemble an anchor set $\mathbb{A}_{3}=\{a_{3}^{(j)}\}$ consists of images of $o_{1}^{E_{1}}$ from $\mathbb{A}_{1}$ and images of $o_{2}^{E_{2}}$ from $\mathbb{A}_{2}$. In the next step, we train the two polices $f^{E_{1}}(g^{E_{1}}_{1}(o^{E_{1}}_{1}),g^{E_{1}}_{2}(o^{E_{1}}_{2}),o^{E_{1}}_{l},\mathbb{A}_{1})$ and $f^{E_{2}}(g^{E_{2}}_{1}(o^{E_{2}}_{1}),g^{E_{2}}_{2}(o^{E_{2}}_{2}),o^{E_{2}}_{l},\mathbb{A}_{2})$ in $E_{1}$ and $E_{2}$ with the Behavior Cloning algorithm Schaal (1996) separately. Then we initialize a stitched policy $f^{E_{1}}(g^{E_{1}}_{1}(o^{E_{1}}_{1}),g^{E_{2}}_{2}(o^{E_{2}}_{2}),o^{E_{1}}_{l},\mathbb{A}_{3})$. This stitched policy is tested in the task $T$ in $E_{3}$. Its performance is measured by its success rate.